OpenStudy (anonymous):

which equation has a slope of -2 and passes through the post (5,0)

OpenStudy (anonymous):

a.y=2x+5

OpenStudy (anonymous):

b.y+2x+10

OpenStudy (raffle_snaffle):

multiple choice eh?

OpenStudy (raffle_snaffle):

well it aint a or b

OpenStudy (zzr0ck3r):

y=mx+b where m = slope and b is the y intercept, hint: (0,b)

OpenStudy (anonymous):

c.y=-2x+5

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(5\quad ,&0) \end{array} \\\quad \\ slope = m= -2 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

The answer is c

OpenStudy (zzr0ck3r):

we should help people get to the answer, not just give them the answer.

OpenStudy (anonymous):

Yeah @zzr0ck3r

OpenStudy (zzr0ck3r):

?